figures/purdue Part 1: Motivation – Constructing aging pathways in yeast Part 2: Tissue Specific Networks and Comparative Analysis

A systems approach to dissecting the tissue-specific architecture of cellular

networks

Ananth Grama

Center for Science of Information Purdue University

San Diego 2014

Work with S. Mohammadi (Purdue), S. Subramaniam (UCSD), and G. Kollias (IBM)

Grama et al.  Tissue-Specific Networks

figures/purdue Part 1: Motivation – Constructing aging pathways in yeast Part 2: Tissue Specific Networks and Comparative Analysis

Outline

1 Part 1: Motivation – Constructing aging pathways in yeast

Overview Materials and Methods

Datasets Tracing Information Flow

Results and Discussion

2 Part 2: Tissue Specific Networks and Comparative Analysis

Motivation Materials and methods

Datasets Algorithmic contributions

Results and discussion

Core alignment graph of housekeeping genes Computing similarity of human tissues with yeast Consistency of alignment p-values Conservation and functional role of tissue selective genes Tissue-selective genes predict tissue-specific pathologies

Grama et al.  Tissue-Specific Networks

figures/purdue

Part 1: Motivation – Constructing aging pathways in yeast Part 2: Tissue Specific Networks and Comparative Analysis  Overview Materials and Methods Results and Discussion

Outline

1 Part 1: Motivation – Constructing aging pathways in yeast

Overview Materials and Methods

Datasets Tracing Information Flow

Results and Discussion

2 Part 2: Tissue Specific Networks and Comparative Analysis

Motivation Materials and methods

Datasets Algorithmic contributions

Results and discussion

Core alignment graph of housekeeping genes Computing similarity of human tissues with yeast Consistency of alignment p-values Conservation and functional role of tissue selective genes Tissue-selective genes predict tissue-specific pathologies

Grama et al.  Tissue-Specific Networks